OpenStudy (anonymous):

How can I simplify this series? sum_{n=0}^{infty}5x(n+r)(n+r-1)c _{n}x ^{n+r-2}+sum_{n=0}^{infty}(7x-6)(n+r)c _{n}x ^{n+r-1}-sum_{n=0}^{infty}(6)c _{n}x ^{n+r} How can I simplify this series? sum_{n=0}^{infty}5x(n+r)(n+r-1)c _{n}x ^{n+r-2}+sum_{n=0}^{infty}(7x-6)(n+r)c _{n}x ^{n+r-1}-sum_{n=0}^{infty}(6)c _{n}x ^{n+r} @Mathematics

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty}5x(n+r)(n+r-1)c _{n}x ^{n+r-2}+\sum_{n=0}^{\infty}(7x-6)(n+r)c _{n}x ^{n+r-1}-\sum_{n=0}^{\infty}(6)c _{n}x ^{n+r}\] Unfortunately, it won't do it symbolically in the first post.

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty}5x(n+r)(n+r-1)c _{n}x ^{n+r-2}\] \[+\sum_{n=0}^{\infty}(7x-6)(n+r)c _{n}x ^{n+r-1}-\sum_{n=0}^{\infty}(6)c _{n}x ^{n+r} \]

OpenStudy (anonymous):

? that's what I put in.

OpenStudy (anonymous):

i know i was just typing it so i could read it. but i have no idea what do to here

OpenStudy (anonymous):

me neither. it's not just for the series, i'm needing to simplify the series to get the answer to a bigger problem

OpenStudy (anonymous):

yeah i figured it was on the way to something

OpenStudy (anonymous):

it's actually to solve a differential equation ;)

OpenStudy (anonymous):

ick

OpenStudy (anonymous):

of second degree

OpenStudy (anonymous):

at least it's homogeneous

OpenStudy (anonymous):

and geometric

OpenStudy (anonymous):

good luck!

OpenStudy (anonymous):

:D thanks